OpenStudy (anonymous):

Find the vertex of the parabola X^2-4x-4y+16=0

OpenStudy (campbell_st):

ok... so rewrite the equation as 4y = (x^2 -4x) + 16 complete the square in x as the 1st step.

OpenStudy (campbell_st):

what do you get...?

OpenStudy (anonymous):

Y=(x^2+x)+4

OpenStudy (campbell_st):

no you need to complete the square so \[4y = (x^2 -4x + 4) + 12\] now factor \[4y = (x -2)^2 + 12\] divide everything by y... \[y^2= \frac{1}{4}(x -2)^2 + 3\] this is one vertex form of the equation the general form is \[y = a(x - h)^2 + k\] where the vertex is at (h, k) match you question to the general form to find the vertex

OpenStudy (anonymous):

So the vertex is (2,3)

OpenStudy (campbell_st):

that's correct.

OpenStudy (anonymous):

Thank you